OpenStudy (anonymous):

Using the chessboard parable, find the total number of grains when all squares up to and including 6 are filled?

OpenStudy (goformit100):

 What is the question It seems unclear to me.

OpenStudy (mathstudent55):

\(1 + 2 \times 1 + 2 \times 2 + 2 \times 4 + 2 \times 8 + 2 \times 16 \) \(= 1 + 2 + 4 + 8 + 16 + 32 \) \(= ~? \)